A 0.106-A current is charging a capacitor that has square plates 4.20 cm on each side. The plate separation is 4.00 mm (a) Find the time rate of change of electric flux between the plates. V.m/s (b) Find the displacement current between the plates. Need Help? İ.eedi. Leater.u
Answer #1

Solution)

Part a) We know, d*phi/dt= i/epsilon not =0.106A/(8.85*10^-12) =1.8*10^9 Vm/s

=========

Part b) displacement current, id= Eo*d*phi/dt

Id= (8.85*10^-12)*(1.8*10^9) = 0.016 A (Ans)

=======

Good luck!:)